# RANGE: Retrieval Augmented Neural Fields for Multi-Resolution Geo-Embeddings

### v1

@@ Abstract
The choice of representation for geographic location significantly impacts the accuracy of models for a broad range of geospatial tasks, including fine-grained species classification, population density estimation, and biome classification. Recent works like SatCLIP and GeoCLIP learn such representations by contrastively aligning geolocation with co-located images. While these methods work exceptionally well, in this paper, we posit that the current training strategies fail to fully capture the important visual features. We provide an information theoretic perspective on why the resulting embeddings from these methods discard crucial visual information that is important for many downstream tasks. To solve this problem, we propose a novel retrieval-augmented strategy called RANGE. We build our method on the intuition that the visual features of a location can be estimated by combining the visual features from multiple similar-looking locations. We evaluate our method across a wide variety of tasks. Our results show that RANGE outperforms the existing state-of-the-art models with significant margins in most tasks. We show gains of up to 13.1% on classification tasks and 0.145 $R^{2}$ on regression tasks. All our code will be released on GitHub. Our models will be released on HuggingFace.
@@ 1 Introduction
To address this issue, we propose a retrieval-augmented strategy that approximates the unique visual information present in a location. There are three key ideas to our solution. First, it is possible to approximate the visual features of a given location by strategically combining the visual features of other similar-looking locations. Second, contrastively trained models like SatCLIP [18] and GeoCLIP [39], while lacking in preserving modality-specific information, are excellent in learning modality alignment [13]. Therefore, it is possible to retrieve semantically aligned images, with alignment scores, for a given location using these models. Finally, we note that the variance of semantic information visible in satellite images across the Earth is relatively low (as opposed to the wider diversity in consumer photographs). Hence, it is possible to capture a large percentage of unique aerial semantics with a limited number of satellite images. By combining these three key ideas, we propose our retrieval-augmented framework, RANGE (Retrieval-Augmented Neural Fields for Multi-Resolution Geo-Embeddings). Our retriever function uses both semantic and spatial alignment to approximate the visual features of a given location using an auxiliary database of image features. Our function is robust to the database size and works well even in limited sample settings. The hyperparameter in the retriever function can be adjusted to generate geo-embeddings at different frequencies. In summary, these are the main contributions of our paper:
@@ 2 Aligning Location and Image is a Multi-View Non-Redundancy Problem
Let (L,H,T) represent location modality, image modality, and geospatial task, respectively. Models like SatCLIP and GeoCLIP align co-located images and locations by minimizing the infoNCE objective. Solving the multimodal infoNCE objective maximizes the lower bound on the mutual information, preserving the shared information within the learned representations [29]. As such, the final learned representations of these models contain only information that is shared across both location and image. This property is useful under the multi-view redundancy assumption which states that the information shared between the two modalities is exactly relevant to solve the underlying task.
In other words, the multi-view redundancy setting assumes that any additional unique information from the image or location would not be useful in solving the task T [37]. However, this assumption does not hold in the location-image learning setting as there is high unique information contained in the image modality that is useful for solving downstream tasks. Formally, we argue that location-image learning falls under the multi-view non-redundancy setting: there is unique information in the image modality that is not shared with the location modality, and that information is relevant for solving our geospatial task T. In this case, contrastive training discards valuable task-relevant information from the image, leading to poor downstream performance [23].
@@
Empirically, there is abundant unique information present in the image features that are relevant for solving downstream tasks. The location representations from contrastively trained models like SatCLIP and GeoCLIP only capture the low-resolution shared information, ignoring all the valuable task-relevant information that is unique to the image. To demonstrate this property concretely, we quantitatively show that there are tasks that can be solved more accurately by adding the image representation to the shared image-location representation from SatCLIP [18]. First, we choose 4 tasks: Biome, Ecoregions, Elevation, and Population Density estimation. We download a corresponding Sentinel-2 image for every data point. For each task, we train one model using the SatCLIP features and another model using the SatCLIP features combined with the SatMAE [6] features of the satellite image [6]. The results in Table 1 show that adding image features improves the performance across all tasks. The increase in performance demonstrates that the images contain unique information relevant to solving the task that is lost during contrastive training. Hence, “location and image” does not follow the assumption of multi-view redundancy.
@@
Although adding corresponding satellite image information to a location embedding improves performance, this is not feasible in practice. For many geospatial tasks, we wish to make predictions across the globe with millions of points. In such cases, using the “image+location” framework would require us to retrieve/store millions of images. We propose an efficient method for approximating the visual information using a compact database. Our results suggest that this approximation can outperform the use of true visual features in some cases.
@@ 4.1 Problem Setup
The training process is shown in Figure 2-(a). This formulation is identical to SatCLIP [18], and so we use the pretrained SatCLIP image and location encoders for simplicity. For query location $g_{i}$, $S_{i}$ encodes the high-resolution visual features since $I$ is a powerful pre-trained image encoder. Both $P_{i}$ and $G_{i}$ encode low-resolution shared information between $g_{i}$ and $s_{i}$. The resulting geo-embedding, $G_{i}$, is sub-optimal as it loses unique visual information crucial for many downstream tasks. To solve this, we propose a retriever function $R_{\tau}$ that allows us to approximate the high-resolution information $S_{i}$ given $g_{i}$. Let $\epsilon$ denote the deviation from the true visual embedding. We can then explicitly add the approximate visual information $\hat{S_{i}}$ to our location embedding $G_{i}$:
@@ 4.2 Retrieval Augmented Neural Fields
type temperature Biome Eco Country Temperature Elevation Population Cali-Housing
@@ 5.1 Quantitative Evaluation on Downstream Tasks
We evaluate the efficacy of RANGE and RANGE+ embeddings on a wide variety of downstream applications. We create our database using the SatCLIP [18] dataset with 82k locations. For RANGE+, we set $\beta$ to $0.5$. We choose 3 classification tasks and 4 regression tasks for this experiment that are used by prior work [18]. The classification tasks are biome classification [8], ecoregion classification [8], and country classification [18]. The regression tasks are air temperature prediction [14], elevation prediction [32], population density prediction [32], and housing price prediction [30]. The objective is to learn a linear predictive model that solves the underlying task. For fairness, we use the same hyperparameters for all tasks, i.e., the same $\tau$ and $\beta$ parameters for each task. The details of the hyperparameters are listed in the supplementary material. We compare our method with state-of-the-art parametric [18, 39, 5, 27, 28] and non-parametric [25, 10] representations. The results of this experiment are shown in Table 2.
The results show that our method outperforms all state-of-the-art models in the classification tasks with significant margins. In biome, ecoregion, and country classification, we advance the state-of-the-art by 13.1%, 4,1%, and 6.4%, respectively. Among the regression tasks, we achieve state-of-the-art performance for elevation and population density prediction with significant gains and narrowly come second in the air temperature prediction task. Our method, however, underperforms in the housing price prediction task. We suspect this is because the visual features that define housing change dramatically over time. The standard California housing dataset is based on a 1990 census, while our retrieved visual features are extracted from 2020 Sentinel data. The results also demonstrate that the hyperparameters for our model are robust and do not need to be tuned for specific tasks to achieve good performance. Table 2 also shows that RANGE+, on average, outperforms RANGE by a small margin, which can be attributed to the spatial smoothness constraint.
@@ 5.4 Qualitative Evaluation of Geo-embeddings
Visualizing the impact of $\beta$ parameter: Secondly, we visualize the impact of the $\beta$ parameter in equation (7). Setting $\beta=1$ gives us the RANGE embeddings, setting $\beta=0$ gives us RANGE-HAVER embeddings, and setting $\beta=0.5$ gives the RANGE+ embeddings. This parameter controls the contribution of spatially-retrieved visual information to the semantically-retrieved visual information. This additional information acts as a spatial smoothness constraint on the RANGE embeddings. At $\beta=0$, we add the maximum constraint, which gives us low-frequency embeddings similar to CSP and SINR. As we increase the value of $\beta$, we increase the frequency of our embeddings, i.e. locations that are spatially close to each other but are semantically different have different representations. Thus, the $\beta$ parameter allows us to explicitly control the smoothness of the RANGE embeddings and, therefore, allows generating location embeddings at multiple frequencies.
@@ 8 Implementation Details
Evaluation on iNaturalist data: For Section 5.2, we conducted a linear probe for each model using the training split of iNaturalist data. However, following prior work [33], we used the “assume negative” loss function, proposed by Cole et al [5]. For RANGE and RANGE+, we use the same hyperparameters as we used for our experiments in Section 5.1. We used the pre-trained “full high-resolution” model by Mac Aodha et al. [25] to get the image-only predictions $P(y\mid I)$ for the iNaturalist test set.

### v2

@@ Abstract
The choice of representation for geographic location significantly impacts the accuracy of models for a broad range of geospatial tasks, including fine-grained species classification, population density estimation, and biome classification. Recent works like SatCLIP and GeoCLIP learn such representations by contrastively aligning geolocation with co-located images. While these methods work exceptionally well, in this paper, we posit that the current training strategies fail to fully capture the important visual features. We provide an information theoretic perspective on why the resulting embeddings from these methods discard crucial visual information that is important for many downstream tasks. To solve this problem, we propose a novel retrieval-augmented strategy called RANGE. We build our method on the intuition that the visual features of a location can be estimated by combining the visual features from multiple similar-looking locations. We evaluate our method across a wide variety of tasks. Our results show that RANGE outperforms the existing state-of-the-art models with significant margins in most tasks. We show gains of up to 13.1% on classification tasks and 0.145 $R^{2}$ on regression tasks. All our code and models will be made available at: https://github.com/mvrl/RANGE.
@@ 1 Introduction
To address this issue, we propose a retrieval-augmented strategy that approximates the unique visual information present in a location. There are three key ideas to our solution. First, it is possible to approximate the visual features of a given location by strategically combining the visual features of other similar-looking locations. Second, contrastively trained models like SatCLIP [18] and GeoCLIP [39], while lacking in preserving modality-specific information, are excellent in learning modality alignment [13]. Therefore, it is possible to retrieve semantically aligned images, with alignment scores, for a given location using these models. Third, we note that the variance of semantic information visible in satellite images across the Earth is relatively low (as opposed to the wider diversity in consumer photographs). Hence, it is possible to capture a large percentage of unique aerial semantics with a limited number of satellite images. By combining these three key ideas, we propose our retrieval-augmented framework, RANGE (Retrieval-Augmented Neural Fields for Multi-Resolution Geo-Embeddings). Our retriever function uses both semantic and spatial alignment to approximate the visual features of a given location using an auxiliary database of image features. Our function is robust to the database size and works well even in limited sample settings. The hyperparameter in the retriever function can be adjusted to generate geo-embeddings at different frequencies. In summary, these are the main contributions of our paper:
@@ 2 Aligning Location and Image is a Multi-View Non-Redundancy Problem
Let (L,H,T) represent location modality, image modality, and geospatial task, respectively. Models like SatCLIP and GeoCLIP align co-located images and locations by minimizing the infoNCE objective. Solving the multimodal infoNCE objective maximizes the lower bound on the mutual information, preserving the shared information within the learned representations [29]. As such, the final learned representations of these models contain only information that is shared across both location and image. This property is useful under the multi-view redundancy assumption which states that the information shared between the two modalities is exactly relevant to solve the underlying task. In other words, the multi-view redundancy setting assumes that any additional unique information from the image or location would not be useful in solving the task T [37]. However, this assumption does not hold in the location-image learning setting as there is high unique information contained in the image modality that is useful for solving downstream tasks. Formally, we argue that location-image learning falls under the multi-view non-redundancy setting: there is unique information in the image modality that is not shared with the location modality, and that information is relevant for solving our geospatial task T. In this case, contrastive training discards valuable task-relevant information from the image, leading to poor downstream performance [23].
@@
Empirically, there is abundant unique information present in the image features that are relevant for solving downstream tasks. The location representations from contrastively trained models like SatCLIP and GeoCLIP only capture the low-resolution shared information, ignoring all the valuable task-relevant information that is unique to the image. To demonstrate this property concretely, we quantitatively show that there are tasks that can be solved more accurately by adding the image representation to the shared image-location representation from SatCLIP [18]. First, we choose 4 tasks: Biome, Ecoregions, Elevation, and Population Density estimation. We download a corresponding Sentinel-2 image for every data point. For each task, we train one model using the SatCLIP features and another model using the SatCLIP features combined with the SatMAE [6] features of the satellite image [6]. The results in Table 1 show that adding image features improves the performance across all tasks. The increase in performance demonstrates that the images contain unique information relevant to solving the task that is lost during contrastive training. Hence, “location and image" does not follow the assumption of multi-view redundancy.
@@
Although adding corresponding satellite image information to a location embedding improves performance, this is not feasible in practice. For many geospatial tasks, we wish to make predictions across the globe with millions of points. In such cases, using the “image+location" framework would require us to retrieve/store/process millions of images. We propose an efficient method for approximating the visual information using a compact database. Our results suggest that this approximation can outperform the use of true visual features in some cases.
@@ 4.1 Problem Setup
The training process is shown in Figure 2-(a). We contrastively align the geolocations and satellite images by minimizing their CLIP objective. This formulation is identical to SatCLIP [18], so we use the pretrained SatCLIP image and location encoders for simplicity. For query location $g_{i}$, $S_{i}$ encodes the high-resolution visual features since $I$ is a powerful pre-trained image encoder. Both $P_{i}$ and $G_{i}$ encode low-resolution shared information between $g_{i}$ and $s_{i}$. The resulting geo-embedding, $G_{i}$, is sub-optimal as it loses unique visual information crucial for many downstream tasks. To solve this, we propose a retriever function $R_{\tau}$ that allows us to approximate the high-resolution information $S_{i}$ given $g_{i}$. Let $\epsilon$ denote the deviation from the true visual embedding. We can then explicitly add the approximate visual information $\hat{S_{i}}$ to our location embedding $G_{i}$:
@@ 4.2 Retrieval Augmented Neural Fields
type temp. Biome Eco Country Temperature Elevation Population Cali-Housing
@@ 5.1 Quantitative Evaluation on Downstream Tasks
We evaluate the efficacy of RANGE and RANGE+ embeddings on a wide variety of downstream applications. We create our database using the SatCLIP [18] dataset with 82k locations. For RANGE+, we set $\beta$ to $0.5$. We choose 3 classification tasks and 4 regression tasks for this experiment as used by prior work [18]. The classification tasks are biome classification [8], ecoregion classification [8], and country classification [18]. The regression tasks are air temperature prediction [14], elevation prediction [32], population density prediction [32], and housing price prediction [30]. Results on climate data from ERA5 are shown in Section 10 of the supplementary material. The objective is to learn a linear predictive model that solves the underlying task. For fairness, we use the same hyperparameters for all tasks, i.e., the same $\tau$ and $\beta$ parameters for each task. The details of the hyperparameters are listed in the supplementary material. We compare our method with state-of-the-art parametric [18, 39, 5, 27, 28] and non-parametric [25, 10] representations. The results of this experiment are shown in Table 2.
The results show that our method outperforms all state-of-the-art models in the classification tasks with significant margins. In biome, ecoregion, and country classification, we advance the state-of-the-art by 13.1%, 4,1%, and 6.4%, respectively. Among the regression tasks, we achieve SoTA performance for elevation and population density prediction with significant gains and narrowly come second in the air temperature prediction task. Our method, however, underperforms in the housing price prediction task. We suspect this is because the visual features that define housing change dramatically over time. The standard California housing dataset is based on a 1990 census, while our retrieved visual features are extracted from 2020 Sentinel data. The results also demonstrate that the hyperparameters for our model are robust and do not need to be tuned for specific tasks to achieve good performance. Furthermore, RANGE+, on average, outperforms RANGE by a small margin, which can be attributed to the spatial smoothness constraint.
@@ 5.4 Qualitative Evaluation of Geo-embeddings
Visualizing the impact of $\beta$ parameter: Secondly, we visualize the impact of the $\beta$ parameter in equation (7). A quantitative evaluation of the $\beta$ parameter is presented in Section 9 of the supplementary material. Setting $\beta=1$ gives us the RANGE embeddings, setting $\beta=0$ gives us RANGE-HAVER embeddings, and setting $\beta=0.5$ gives the RANGE+ embeddings. This parameter controls the contribution of spatially-retrieved visual information to the semantically-retrieved visual information. This additional information acts as a spatial smoothness constraint on the RANGE embeddings. At $\beta=0$, we add the maximum constraint, which gives us low-frequency embeddings similar to CSP and SINR. As we increase the value of $\beta$, we increase the frequency of our embeddings, i.e. locations that are spatially close to each other but are semantically different have different representations. Thus, the $\beta$ parameter allows us to explicitly control the smoothness of the RANGE embeddings and, therefore, allows generating location embeddings at multiple frequencies.
@@ 8 Implementation Details
Evaluation on iNaturalist data: For Section 5.2, we conducted a linear probe for each model using the training split of iNaturalist data. However, following prior work [33], we used the “assume negative" loss function, proposed by Cole et al [5]. For RANGE and RANGE+, we use the same hyperparameters as we used for our experiments in Section 5.1. We used the pre-trained “full high-resolution" model by Mac Aodha et al. [25] to get the image-only predictions $P(y\mid I)$ for the iNaturalist test set.
@@ 11 Visualizing Geo-Embeddings at Country Scale
12 Geoprior Evaluation with Training-free Baselines
In Section 5.2, we evaluated different training-based location encoding methods on geoprior task using iNaturalist data. Here, we show the results of using training-free location encoding methods. Table 7 shows that RANGE models outperform the training-free baselines.
top-1 top-3 top-5 top-10
Direct 63.5 81.7 87.0 91.8
Cartesian 65.0 82.7 87.7 92.3
Wrap 65.4 83.1 87.9 92.5
SphereC+ 69.5 85.5 89.9 93.5
SphereM+ 70.8 86.3 90.5 93.9
RANGE 75.2 89.6 92.9 95.5
RANGE+ 75.1 89.5 92.8 95.5
Table 7: Top-k classification accuracy on INat-2018 test split: Comparing our model with training free baselines.
13 Computational Cost
The retrieval process incurs some added computational cost. However, our setup makes this process highly efficient. Generating the feature bank is a one-time operation, which is inexpensive for a few thousand images (Section 5.3). Second, the retrieval process is completely vectorized and highly efficient. For reference, when using the 10k database, computing the RANGE embeddings for 1 million input locations takes less than 65 seconds on our CPU and less than 10 seconds on our H100 GPU, making our method efficient for any practical usage.